# A Parametric Study on the Pulsation of Burner-Stabilized CH4-O2 Flames Moderated by CO2 Addition – Frequencies, Modes and Regime Diagrams

### v2

@@ Abstract
This study investigated the pulsating instability of burner-stabilized premixed CH4-O2 flames at various levels of CO2 dilution. Experiments were conducted using a water-cooled porous-plug burner of 18 mm diameter over a wide range of mixture compositions and flow rates, during which time-resolved measurements of flame chemiluminescence and gas temperature were obtained. The primary oscillation frequencies of the pulsating flames were determined using fast Fourier transform and harmonic power analysis. Phase-locked analysis of the chemiluminescence images revealed an interesting mode-transition phenomenon of the flame oscillations. Under fuel-rich conditions with relatively low heat release rates and low flow rates, the flames exhibited quasi-periodic single-mode oscillations. These oscillations were modulated by low-frequency flame flickering instabilities at elevated flow rates, which created sidebands around the primary oscillation frequency. At higher heat release rates, the flickering instability further triggered mode splitting, eventually leading to multi-mode oscillations. Regime diagrams of the flame oscillation modes, as well as the stability boundaries, were obtained under various fuel flow rates. These findings can be useful for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
@@
This work presents, to the authors’ knowledge, the first systematic measurements of pulsating flame instabilities in burner-stabilized premixed methane-oxygen flames across a wide range of experimental conditions (equivalence ratios, dilution factors and flow rates). The combination of a porous-plug burner and oxygen-enriched combustion amplifies weak phenomena that commonly occur under extreme conditions of very high Lewis number, making them observable under moderate conditions that are easily accessible in a laboratory. Additionally, this work reports direct measurements of the spatiotemporal evolution of flame chemiluminescence and gas temperature during pulsation cycles. Based on phase-locked analysis of the data, this study has revealed an interesting mode-transition phenomenon of the flame pulsation. The results of this work should prove useful for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
@@ 1 Introduction
An important type of such instability is the thermo–diffusive instability that arises from the imbalance between heat and mass transfer [7, 8, 9]. Depending on the effective Lewis number (Le) of the reactive mixture, the flame can exhibit either spatial cellular structures (at Le $<$ 1) or temporal pulsation of the reaction front (at Le $>$ 1). Although the former has been extensively investigated, especially in lean combustion systems of lightweight fuels such as hydrogen and methane [10, 11, 12, 13, 14, 15, 16, 17, 18], experimental studies on the latter remain relatively scarce.
@@ 2.2 Frequency analysis and phase-locked cycle averaging of chemiluminescence signal
A harmonic power analysis method was employed to extract the fundamental frequency of oscillation, $f_{osc}$, by exploiting the fact that the frequency content of a periodic physical event consists of harmonics that are phase-locked, whereas the frequency content of noise or random fluctuations is uncorrelated. Specifically, $f_{osc}$ was determined by solving a constrained optimization problem, namely maximizing the total power of harmonics within a search window $[f_{min},f_{max}]$ estimated from the NPSD result, as shown below.
@@ 2.3 Axisymmetric reconstruction of temperature fields from laser absorption measurement
From the time sequence of $\alpha(t;y_{i})$, an absorbance spectrum $\alpha(\nu;y_{i})$ was obtained based on the laser wavenumber modulation profile $\nu(t)$ characterized prior to the flame experiment. A total of 12 absorption features in the spectral region of 2392.5 to 2397 $\rm cm^{-1}$ were analyzed, and the integrated absorbance over each selected transition feature (denoted by $[\nu_{k-},\nu_{k+}]$) was used in the axisymmetric reconstruction, i.e.,
[…]

### v3

# Pulsation of Burner-Stabilized CH4-O2 Flames Moderated by CO2 Addition – Frequencies, Modes and Regime Diagrams

@@ Abstract
This study investigated the pulsating instability of burner-stabilized premixed CH4-O2 flames at various levels of CO2 dilution. Experiments were conducted using a water-cooled porous-plug burner of 18 mm diameter over a wide range of mixture compositions and flow rates, during which time-resolved measurements of flame chemiluminescence and gas temperature were obtained. The primary oscillation frequencies of the pulsating flames were determined using fast Fourier transform and harmonic power analysis. Phase-locked analysis of the chemiluminescence images revealed an interesting mode-transition phenomenon of the flame oscillations. Under fuel-rich conditions with relatively low heat release rates and low flow rates, the flames exhibited quasi-periodic single-mode oscillations. At elevated flow rates, these oscillations were modulated by low-frequency flame flickering instabilities, which created sidebands around the primary oscillation frequency. At higher heat release rates, the flickering instability further triggered mode splitting, eventually leading to multi-mode oscillations. Regime diagrams of the flame oscillation modes, as well as the stability boundaries, were obtained under various fuel flow rates. These findings can be useful for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
@@
This work presents, to the authors’ knowledge, the first systematic measurements of pulsating flame instabilities in burner-stabilized premixed methane–oxygen flames. Utilizing a custom-built miniature porous-plug burner and oxygen-enriched combustion, we obseverd the intrinsic flame pulsation typically occurring at very high Lewis numbers under moderate laboratory conditions. The spatiotemporal evolution of flame chemiluminescence and gas temperature was resolved throughout individual pulsation cycles, and a phase-locked analysis revealed an interesting mode-transition phenomenon of flame pulsation. The results of this work are useful for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
@@ 1 Introduction
An important type of such instability is the thermo–diffusive instability that arises from the imbalance between heat and mass transfer [7, 8, 9]. Depending on the effective Lewis number (Le) of the reactive mixture, the flame can exhibit either spatial cellular structures (at Le << 1) or temporal pulsation of the reaction front (at Le >> 1). Although the former has been extensively investigated, especially in lean combustion systems of lightweight fuels such as hydrogen and methane [10, 11, 12, 13, 14, 15, 16, 17, 18], experimental studies on the latter remain relatively scarce.
@@ 2.2 Frequency analysis and phase-locked cycle averaging of chemiluminescence signal
A harmonic power analysis method was employed to extract the fundamental frequency of oscillation, $f_{osc}$, by exploiting the fact that the frequency content of a periodic physical event consists of harmonics that are phase-locked, whereas the frequency content of noise or random fluctuations is uncorrelated. Specifically, $f_{osc}$ was determined by solving a constrained optimization problem, namely maximizing the total power of harmonics within a search window $f_{min},f_{max}]$ estimated from the NPSD result, as shown below.
@@ 2.3 Axisymmetric reconstruction of temperature fields from laser absorption measurement
From the time sequence of $\alpha(t;y_{i})$, an absorbance spectrum $\alpha(\nu;y_{i})$ was obtained based on the laser wavenumber modulation profile $\nu(t)$ characterized prior to the flame experiment. A total of 12 absorption features in the spectral region of 2392.5 to 2397 $\rm cm^{-1}$ were analyzed, and the integrated absorbance over each selected transition feature (denoted by $\nu_{k-},\nu_{k+}]$) was used in the axisymmetric reconstruction, i.e.,
$A_{k}(y_{i})=\int_{\nu_{k-}}^{\nu_{k+}}\alpha(\nu;y_{i})d\nu$ (5)